OpenStudy (anonymous):

Ok I am trying to solve this equation x^2+12x+4=0 using the quadratic equation. I have gotten to -12+/-sqrt(128). What do I do after this, how do I simplify the 128?

OpenStudy (mertsj):

\[\sqrt{128}=\sqrt{64(2)}=8\sqrt{2}\]

OpenStudy (anonymous):

Ok thank you so much so my answer would be -6+/-8sqrt(2)

OpenStudy (mertsj):

What about the denominator?

OpenStudy (anonymous):

/2

OpenStudy (mertsj):

So then, you have to divide out the two

jimthompson5910 (jim_thompson5910):

you are close, the answer is actually -6+/-4sqrt(2)

OpenStudy (mertsj):